OpenStudy (anonymous):

Find all solutions to the equation: 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (cwrw238):

7 sin^2x - 14 sin x + 7 = 0 first we need to solve this for sin x

OpenStudy (anonymous):

ok. do we need to isolate the sinx from the original equation or is there another method for doing this?

OpenStudy (anonymous):

then factor it 7(sin^2x-2sinx+1)=0

OpenStudy (anonymous):

now do basic operations with it sin^2x-2sinx+1=0 (sinx-1)^2=0

OpenStudy (anonymous):

then sinx=1

OpenStudy (anonymous):

and solve the last x=pi/2 or 90^deg

OpenStudy (anonymous):

Oh, so you got rid of the 7 attached outside of the factored equation, correct?

OpenStudy (anonymous):

uyes

OpenStudy (anonymous):

Ok, I understand... I'm still getting the hang of these kind of problems. Thank you very much!

OpenStudy (anonymous):

cool then